OpenStudy (anonymous):

Sqrt(x-95)+8=15 SOLVE FOR X Help me please :)

OpenStudy (matt101):

\[\sqrt{x-95}+8=15\]\[\sqrt{x-95}=7\]\[x-95=49\]\[x=144\]

OpenStudy (anonymous):

Quite so my good man! QUITE SO!!

OpenStudy (anonymous):

Awesome, thank you :)